10 moles of an ideal gas expands irreversibly against an unknown constant external pres- sure, Pert, from an initial volume Vİ-1 L to a final volume ½ 11 L. In the process, the temperature of the gas falls from T350 K to T2 250 K, and it absorbs heat q+7 L atm from the surroundings. (a) What is the external pressure, Pert (in atm)? [Note: this is an ideal gas, so its internal energy depends only on its temperature.] (b) What is the final pressure of the gas, P2 (in atm)?
Answer #1

a) using the universal gas equation we have

PV = nRT

P*1 = 10 * 0.0821*350

P = 287.35 atm

(b) P1V1/T1 = P2V2/T2

287.35*1 / 350 = P2*11 / 250

P2 = 18.65 atm